OpenStudy (anonymous):

Given the function f(x)=2/x, calculate the difference quotient f(x)−f(a) / x−a

OpenStudy (cwrw238):

(2/x - 2/a) / (x-a) = 2a - 2x 1 ------ * ------- ax (x-a) = -2 (x - a) * 1 --------- --- ax (x-a) = -2 -- ax